# Entanglement detection with quantum support vector machine (QSVM) on near-term quantum devices

### v1

@@ 2 Quantum Support Vector Machines
where $\sigma$ represents one of the Pauli matrices, and $W_{k}$ represents a two-qubit gate. The parameters $\boldsymbol{\vec{\theta}}=(\theta_{1},\theta_{2},...)$ determine the transitions between states, enabling the PQC to represent the classical decision function in quantum form. The loss function depends on the VQC technique (ansatz) and consists of a series of quantum gates and operations. The optimization parameters $\mu_{i}(\boldsymbol{\vec{\theta}})$ are derived from the ansatz and are equivalent to the squared magnitude of the expectation value of the operator $U(\boldsymbol{\vec{\theta}})$ applied to the initial state. As outlined in Appendix A Eq.27, we adopt the classical SVM loss function method, enhanced by the integration of quantum-based parameters for improved performance.The quantum-based loss function is expressed as:
@@
This framework allows for the efficient detection of entangled and separable quantum states, paving the way for broader applications of QSVMs in quantum machine learning  1.
@@ 2.1.2 Entangled states
Concurrence can be described as an informative item in quantum circuits for labeling data and its value in this circuit is equivalent to $c=\sin(\theta_{0})\sin\left(\frac{\theta_{1}}{2}\right)$ where $\sin(\theta_{0})$ is the control qubits degree of initial superposition (Fig. 6).
@@ 2.2 Entangling Capability and Expressibility
According to Figure 8, it can be seen that in the one blocks, we need a smaller amount of entanglement for the quantum circuit in order to reach a high accuracy value. In this case, it shows that the increase of blocks will be accompanied by an increase in cost and space for the quantum circuit, which in this figure shows that this increase in the block and volume of the quantum circuit will not necessarily increase the amount of accuracy in quantum processing. The values of accuracy and entanglement capability in each circuit is shown in Tab. 1
@@
Here, $D_{k}$ is the Kullback-Leibler divergence and $P_{PQC}(F,\theta)$ represents the estimated probability distribution of fidelities obtained by sampling states generated from a PQC. For an ensemble of Haar-random states, the probability density function of fidelities is given by $P_{Haar}(F)=(N-1)(1-F)^{(N-2)}$, where F and N are the fidelity and dimension of the Hilbert space, respectively. Expressibility measures how well the state distribution from a PQC matches that of Haar random states, indicating the extent to which the states span the Hilbert space. High expressibility or low Kullback-Leibler divergence values indicate how effectively the PQC can generate all possible quantum states within the Hilbert space (For further details and an example involving a single-qubit state, see Ref. [53]). Figure 9 illustrates that as the number of blocks in a quantum circuit increases, expressibility also rises. This means the quantum states generated by the circuits cover a broader region of the Hilbert space. Consequently, the accuracy of QSVM in detecting quantum states improves with increased expressibility. Also, the values of accuracy and expressibility in each circuit is shown in Tab. 2
@@
Table 1: Comparison of Accuracy and Entangling Capability for Different Circuits.
@@
Table 2: Comparison of Accuracy and Expressibility for Different Circuits.
@@ 3 Investigating IBMQ through Experimental Trials
Our method necessitated classical optimization, so we incorporated session definitions in our code. The number of circuit shots was set to 8192. During training data sampling, the data dimension was reduced to $n=2$ with a size of $N=4$ to mitigate decoherence effects.Twenty test datasets were generated with random $\theta$ values using the circuit shown in Fig. 6. The Simultaneous Perturbation Stochastic Approximation (SPSA) minimization tool was chosen as it aligns well with the random nature of quantum states and the data sampling [55, 56].
The elapsed time for each circuit’s runtime process is presented in table 3.

### v2

@@ 2 Quantum Support Vector Machines
where $\sigma$ represents one of the Pauli matrices, and $W_{k}$ represents a two-qubit gate. The parameters $\boldsymbol{\vec{\theta}}=(\theta_{1},\theta_{2},...)$ determine the transitions between states, enabling the PQC to represent the classical decision function in quantum form. The loss function depends on the VQC technique (ansatz) and consists of a series of quantum gates and operations ( see Figuers of .4 and 5 ). The optimization parameters $\mu_{i}(\boldsymbol{\vec{\theta}})$ are derived from the ansatz and are equivalent to the squared magnitude of the expectation value of the operator $U(\boldsymbol{\vec{\theta}})$ applied to the initial state. As outlined in Appendix A Eq.27, we adopt the classical SVM loss function method, enhanced by the integration of quantum-based parameters for improved performance.The quantum-based loss function is expressed as:
@@
This framework allows for the efficient detection of entangled and separable quantum states, paving the way for broader applications of QSVMs in quantum machine learning(see Fig.  1 ).
@@ 2.1.2 Entangled states
Concurrence can be described as an informative item in quantum circuits for labeling data and its value in this circuit is equivalent to $c=\sin(\theta_{0})\sin\left(\frac{\theta_{1}}{2}\right)$ where $\sin(\theta_{0})$ is the control qubits degree of initial superposition (see Fig. 6).
@@ 2.2 Entangling Capability and Expressibility
According to Figure 8, it can be seen that in the one blocks, we need a smaller amount of entanglement for the quantum circuit in order to reach a high accuracy value. In this case, it shows that the increase of blocks will be accompanied by an increase in cost and space for the quantum circuit, which in this figure shows that this increase in the block and volume of the quantum circuit will not necessarily increase the amount of accuracy in quantum processing. The accuracy and entanglement capability values for circuits C1 through C19 (see Fig.11) are provided in Tab. 1
@@
Here, $D_{k}$ is the Kullback-Leibler divergence and $P_{PQC}(F,\theta)$ represents the estimated probability distribution of fidelities obtained by sampling states generated from a PQC. For an ensemble of Haar-random states, the probability density function of fidelities is given by $P_{Haar}(F)=(N-1)(1-F)^{(N-2)}$, where F and N are the fidelity and dimension of the Hilbert space, respectively. Expressibility measures how well the state distribution from a PQC matches that of Haar random states, indicating the extent to which the states span the Hilbert space. High expressibility or low Kullback-Leibler divergence values indicate how effectively the PQC can generate all possible quantum states within the Hilbert space (For further details and an example involving a single-qubit state, see Ref. [53]). Figure 9 illustrates that as the number of blocks in a quantum circuit increases, expressibility also rises. This means the quantum states generated by the circuits cover a broader region of the Hilbert space. Consequently, the accuracy of QSVM in detecting quantum states improves with increased expressibility. Also, the values of accuracy and expressibility for circuits C1 through C19 (see Fig.11) are provided in Tab. 2
@@
Table 1: We utilized various circuits, ranging from C1 to C19, as detailed in the Appendix C, and conducted a comparison of their Accuracy and Entangling Capability.
@@
Table 2: We utilized various circuits, ranging from C1 to C19, as detailed in the Appendix C, and conducted a comparison of their Accuracy and Expressibility.
@@ 3 Investigating IBMQ through Experimental Trials
Our method necessitated classical optimization, so we incorporated session definitions in our code. The number of circuit shots was set to 8192. During training data sampling, the data dimension was reduced to $n=2$ with a size of $N=4$ to mitigate decoherence effects.Twenty test datasets were generated with random $\theta$ values using the circuit shown in Fig. 6. The Simultaneous Perturbation Stochastic Approximation (SPSA) minimization tool was chosen as it aligns well with the random nature of quantum states and the data sampling  [55, 56]. The optimization steps up to the minimum stage with 200 iterations for Table  4 are shown in Figure  10. The elapsed time for each circuit’s runtime process is presented in table 3.
@@ B.3 Software
B.4 Data availability
The numerical data generated in this work are available from the corresponding author on reasonable request.